OpenStudy (anonymous):

Solve each quadratic equation by Factoring x^2 - 15x + 56 = 0 2x^2 + 4x - 70 + 0 2z^2 + 8x = 5x + 20 x^2 + 6x + 9 = 0 please show how you received the answer by factoring thank you.

OpenStudy (anonymous):

(x-8)(x-7)=0, number 1 -8*-7=56, (2x-10)(x+7)=0, number 2 -10*7=-70 2x^2+3x-20=0, (2x-5)(x+4)=0, number 3 (x+3)^2=0, number 4

OpenStudy (anonymous):

oh crap the last one was supposed to be x^2 + 6x + 9 = 9, i dont understand that

OpenStudy (anonymous):

and how is 2x^2 distributed??

OpenStudy (anonymous):

the last one will be if it is x^@+6x+9=9, subtract the 9 to get the one side = 0 x^2+6x=0 GCF : x x(x+6)=0